if y=x/x+2 prove that xdy/dx=(1-y)y

Grade:12

2 Answers

bharat bajaj IIT Delhi
9 years ago
dy/dx = 2/(x+2)^2

Lhs = 2x/(x+2)^2
Rhs= (1-y)y = 2/(x+2). x/(x+2)
Rhs = 2x/(x+2)^2

Lhs =rhs

Thanks
Bharat Bajaj
Askiitians faculty
Arun Kumar IIT Delhi
9 years ago
Hi

Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty